Michelle and Thea are running in the Fun Run. Michelle runs at 7 kilometers per hour and Thea at 5 kilometer per hour. If they start at the same time, how long will it be before Michelle is one-fourth kilometer ahead of Thea?

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16946)

Michelle is 2 km ahaed every 2 hours
let her be 1/4 km after x hours
x= (1/4)/2
x=1/8 hour =7 .5 minutes

Answer by ikleyn(52781)

.

It was solved today under this link

https://www.algebra.com/algebra/homework/word/travel/Travel_Word_Problems.faq.question.1189809.html